Find three consecutive negative integers such that the sum of the greatest and twice the smallest is equal to four times the middle integer. So far 4(n-1)=n+2(n-2) but not ending up with three consecutive negative integers. Thank you
Answer by ikleyn(52781)

.

The numbers are x, x+1 and x+2.


The equation is  


    (x+2) + 2x = 4*(x+1).


Simplify and find x


    x + 2 + 2x = 4x + 4

    3x         = 4x + 2

    -x = 2.

     x = -2.


The numbers are -2, -1, 0.


CNECK.  a)  0 + 2*(-2) = -4.  b)  4*(-1) = -4.

        So,  a) = b).   ! Correct !


But "0" is not negative, so the condition is not precisely correct.

Answered.